# Modeling of AR 12760 with GX Simulator and Evidence for the Extended Transition Region in Peripheral Active Region Loops

### v1

@@ 3.3 Fits to other AIA bands
4 Discussion
4.1 Heating Parameters
@@
Figure 14: $\sum\delta^{2}$ values for the fit to the (a) 211 Å data and (b) 335 Å data using a 8″ resolution map for the data to model comparison. The red asterisk shows the best fit location, and the dashed line shows a fit to the minima in each column. The color bar shows the $\sum\delta^{2}$ values normalized to the minimum $\sum\delta^{2}$.
Our most reliable fits are the low resolution (8″) fits to the 211 and 335 Å data, because the lower resolution smooths out some of the inaccuracies in the magnetic field model. Figure 14 shows plots of our model-to-data comparison metric, $\sum\delta^{2}$, as a function of $a$ and $b$, where each fit has been optimized for $q_{0}$ for that fitted $a,b$ pair (see Figure 18 and 20). The best fit for 211 Å, shown by the asterisk, is $a=1.5$, $b=1$, and $q_{0}=7\times 10^{-3}$ erg cm-3 s-1. If we consider the $q_{0}$ values for fits within 20% of the minimum $\sum\delta^{2}$ value then we get $q_{0}$ in the range $3\times 10^{-3}$ and $11\times 10^{-3}$ erg cm-3 s-1. For 335 Å it is $a=0.5$, $b=1.5$, and $q_{0}=4\times 10^{-3}$ erg cm-3 s-1. The fits within 20% of the minimum include $q_{0}$ values from $3\times 10^{-3}$ to $7\times 10^{-3}$ erg cm-3 s-1.
@@
Table 1: 211 and 355 Å best fit comparisons
A recurring feature in these plots of model-to-data comparison metrics is the clear diagonal feature ranging from low $a$, high $b$ values on the top-left to high $a$ to low $b$ values towards the bottom-right. This feature is consistent with other studies with EUV (Ugarte-Urra et al., 2019) and microwave data (Nita et al., 2023; Fleishman et al., 2025; Kuznetsov et al., 2025). For the 8″ resolution 211 Å fit we found that this feature roughly corresponds to $b\approx 1.6-0.6a$ and for the 335 Å fits we get $b\approx 1.7-0.6a$. That our best fits for each wave band do not fall exactly on these lines is indicative of the coarseness of our ($a,b$) grid and uncertainties in our fitting of diagonal features. More specifically, if we assume a 0.25 uncertainty (half a ($a,b$) grid step) on both $a$ and $b$ we have $b$ intercepts of $1.61\pm 0.12$ and $1.70\pm 0.13$ and slopes of $0.56\pm 0.07$ and $0.59\pm 0.08$ for 211 Å and 335 Å, respectively, so the results for the two bands are the same within uncertainties.
We note that although the best $(a,b)$ values are different for the two wave bands, they both fall close to these best-fit lines. Table 1 compares the best model-to-data comparisons of the two wave bands. We see that for both wave bands the best ($a,b$) values fall within the 10% of minimum $\sum\delta^{2}$ zone for ($a,b$) values of the other waveband. This gives an idea, albeit somewhat qualitative, of our confidence in our fit values. The model-to-data comparisons for the other bands are sufficiently poor (see Sections 3.2 and 3.3) that we do not include them in this analysis.
@@
$B\propto L^{-\gamma}$ (2)
(Klimchuk & Porter, 1995) then we have $\langle Q\rangle\propto L^{-\gamma a-b}$. Model fits should be similarly good when $\gamma a+b=$ constant; call it $b_{0}$. Good fits should therefore fall along a straight line in a plot of goodness of fit in the $(a,b)$ plane:
@@
Our results shown in Figure 14 suggest that $b_{0}\approx 1.6$ or 1.7 and $\gamma\approx 0.6$. As stated above, other analyses also show that similarly good fits fall along a straight line. Fleishman et al. (2025) using GX Simulator to model optically thick gyroresonant microwave emission observed by the Siberian Solar Radio Telescope (SSRT) and Nobeyama Radioheliograph (NoRH) found b=2.41.0ab=2.4-1.0a at 5.7 GHz and b=2.20.7ab=2.2-0.7a at 17 GHz and a best solution at ($a,b$) = (0.6, 1.8). Kuznetsov et al. (2025) analyzed eight active regions in 6, 8, and 10 GHz with the Siberian Radioheliograph and found that $b_{0}$ was mostly between 1.6 and 3.2, with the distribution peaking at 2.5, and $\gamma$ ranged from 0.2 to 1.2, with a most probable value of 0.55. It is also interesting that Porter & Klimchuk (1995) concluded that $\langle Q\rangle\propto L^{-2}$ from a study of soft X-ray loops observed by Yohkoh, which can be compared to the $b_{0}$ values derived from GX Simulator.
@@
With this in mind, we fit the $B_{\mbox{\scriptsize avg}}$ vs $L$ dependence in our Magnetic Model 1, as shown in Figure 15, using a least squares fit in log-log space. We derived $\gamma=0.8$. The plot also shows a line with a slope corresponding to $\gamma=0.6$ as suggested by the GX Simulator modeling. The value derived directly from the magnetic model is slightly higher than the value derived from $\sum\delta^{2}$ plots, but it is clear from the distribution of points in Fig. 15 that the distribution of the field lines in the model is more complex that a simple power-law, with noticeable substructure. We expect Equation 2 to be a general trend rather than an equation applying precisely to all parts of the active region. These results are, however, consistent with the idea that the diagonal feature in GX Simulator model-to-data comparison metrics is due to a dependence of $B_{\mbox{\scriptsize avg}}$ on $L$.
@@
Figure 15: $B_{\mbox{\scriptsize avg}}$ vs $L$ for 50,000 randomly selected points in Magnetic Model 1. Lines shown are an Ordinary Least Squares (OLS) regression to the entire set of points (solid blue line) and a line with a slope of $\gamma=0.6$, as derived from the GX Simulator fits (red dashed line). This line is forced to go through the center of mass of the magnetic model points (red circle).
@@
We should note that Fleishman et al. (2025) and Kuznetsov et al. (2025) are looking at different active regions, which in addition to a difference in wavebands and a different model-to-data comparison metric could lead to differences in the properties of the zones of best fit. Kuznetsov et al. (2025) in particular finds a range of values over the many active regions they modeled.
@@
However, the broad agreement of these results is very encouraging. This validates the power-law dependencies assumed in Equations 1 and 2. Had similarly good fits not fallen along a straight line, one of more of the dependencies would be invalidated.
4.2 Transition Region Emission
@@
An important issue in our modeling effort is that the emission from long features at the periphery of the active region is missing from the models. For this active region these are dominant features in the cooler bands, 131, and 171 Å, and significant contributors to other bands with peak contributions below about 1.6 MK (94 and 193 Å). We think this is due to an unrealistic distribution of the transition region emission in these long loops. In this work and other modeling studies, the transition region is defined to be the locations in the loops where thermal conduction serves as a heating term rather than a cooling term (Vesecky et al., 1979). This means that the emission does not have to be confined to a particular temperature or altitude range, as it is often defined observationally. In general, the transition region extends from the top of the chromosphere to about 10% of a loop half-length ($0.05L$) (see Klimchuk et al., 2008; Cargill et al., 2022). An example of this is shown in Figure 16, which depicts the temperature as a function of position of a 1D hydrodynamic model of a loop in equilibrium with steady uniform heating. The model was computed with the ARGOS code (Antiochos et al., 1999) and is broadly representative of static equilibrium loops with a wide range of lengths and heating rates. The stars show the top of the transition region at $0.054L$ from the loop footpoints. This same spatial fraction applies approximately to other loops, but the temperature of the top of the transition region varies, always about 60% of the apex temperature.
@@
Figure 16: The temperature of a equilibrium loop with uniform heating as modeled by ARGOS. The stars show the location of the top of the transition region/base of the corona. It is about 5.4% of full loop length, $L$, from the top of the chromosphere
@@
For relatively short loops it is sufficient to assume that the transition region emission would be confined to a limited range approximated by a voxel in our model cube. However, based on Magnetic Model 2, the loops at the periphery of the active region likely have L=200700L=200-700 Mm, implying transition regions on the order of 10-50″ in the AIA images, much larger than the pixels immediately around the loop base and roughly consistent with the large cool loops we see in the data. Thus, in order to get a better fit to EUV images showing extended emission in loop legs, it will be necessary to account for the finite spatial extent of the transition region. Note that it is only the upper transition region, the upper third as determined by temperature, where this is important. The temperature scale lengths in the lower and even middle transition region are extremely short. The amount of transition region emission predicted by EBTEL is reasonable (although see caveats about general transition region modeling below), but because it is confined to the footpoints of the model loops, the footpoints are too bright and the extended legs do not appear in model images.
@@
This explanation would explain Feature A, the fan-loop feature in Figure 10. It is not as clear if it accounts for Feature B in that plot, in which the emission appears to extend farther than $0.05L$ along a magnetic tube.
@@
As discussed in Fleishman et al. (2025), one possible way to obtain more information about the transition region is to combine analysis of microwave emission, which is sensitive to only the corona, with EUV emission, which includes both transition region and coronal components. This will help to separate out the emission from the two sources. Another approach would be to consider active regions from multiple points of view using AIA combined with either STEREO/EUVI (Solar TErrestrial RElations Observatory/Extreme UltraViolet Imager; Howard et al., 2008) or Solar Orbiter/EUI (Extreme Ultraviolet Imager; Rochus et al., 2020) near quadrature, so that lower altitude transition region emission could be separated out. However, for both of these approaches it would be important to consider the extended transition region emission from long loops discussed in this paper.
In addition to implementing a more accurate distribution of transition region emission in loops, there are other ways in which the transition region modeling could be improved. One important advance would be to include loop expansion in the model. Both EBTEL and the multi-strand MHD simulations have this capability. This is a significant undertaking as it adds a third dimension to the model lookup table. The DEM would then depend on the expansion factor, $\Gamma$, in addition to $\langle Q\rangle$ and $L$. A more simple approach, which has very recently been added to GX Simulator would be to add a multiplicative factor accounting for expansion to the formula for $\langle Q\rangle$ (Equation 1). How this affects model intensities should be investigated.
4.3 Further Considerations
@@
In addition to the treatment of the transition region there are other areas in which our modeling could be improved. As stated above, the details of the core magnetic field lines do not reflect the data well. This is likely because this active region had a rather weak magnetic field. It was selected because we wished to have a very quiet active region so as to avoid solar flares in the time-average images. Also, we wished to avoid sunspots because it is highly unlikely that sunspots obey the same heating scaling law that applies throughout most of the active region. We preferred to avoid this complication in our initial study, but of course sunspots must be included in an eventual complete model. A more typical active region, with stronger fields, may be easier to model magnetically. It would optimally make sense to have an iterative fit in which both the plasma and magnetic field parameters could be varied (see Fleishman et al. (2021a) for a case where fits to radio data were done using multiple magnetic models), but that is beyond the scope of this paper and our readily available computing power.
@@
Another area to consider is the assumptions going into the heating calculations in the lookup tables. Fleishman et al. (2025) and Kuznetsov et al. (2025) considered the various lookup tables currently available in GX Simulator and, in analyses utilizing microwaves, concluded that the best fits for the diffuse coronal component come from assuming heating parameters based on higher frequency nanoflares than we used for this study or even constant heating. So, ultimately, heating with various nanoflare distributions should be considered for the EUV as well. The power-law distribution of heating events used here is a common one in the field, but our multi-strand MHD simulations indicate that the heating per unit area input to a magnetic strand, the quantity used by EBTEL, obeys a log-normal distribution (Sow Mondal et al., 2025). Furthermore, the delay and intensity of the nanoflares are uncorrelated, contrary to the assumption in the current GX Simulator lookup tables. These properties must be included when new lookup tables are computed.
@@
Although we concluded that it would not have helped in our case, it would be good to develop a way to estimate the loop length of field lines that exit the model box. At the moment the options are to either ignore the plasma in such lines or assume it is hydrostatic, neither of which is consistent with the modeling of the loops that close in the box. The calculation and manipulation of large volume NLFFF models needed to include more loops within the modeled box uses a lot of computer resources, so a way of estimating $L$ and $B_{\mbox{\scriptsize avg}}$ for large loops exiting smaller boxes would be helpful.
5 Summary and Conclusions

### v2

@@ III.3 Fits to other AIA bands
IV Discussion
IV.1 Heating Parameters
@@
Figure 15: $\sum\delta^{2}$ values for the fit to the (a) 211 Å data and (b) 335 Å data using a 8″ resolution map for the data to model comparison. The red asterisk shows the best-fit location, and the dashed line shows a fit to the minima in each column. The color bar shows the $\sum\delta^{2}$ values normalized to the minimum $\sum\delta^{2}$.
Our most reliable fits are the low-resolution (8″) fits to the 211 and 335 Å data, because the lower resolution smooths out some of the inaccuracies in the magnetic field model. Figure 15 shows plots of our model-to-data comparison metric, $\sum\delta^{2}$, as a function of $a$ and $b$, where each fit has been optimized for $q_{0}$ for that fitted $a,b$ pair (see Figures 18 and 21). The best fit for 211 Å, shown by the asterisk, is $a=1.5$, $b=1$, and $q_{0}=7\times 10^{-3}$ erg cm-3 s-1. If we consider the $q_{0}$ values for fits within 20% of the minimum $\sum\delta^{2}$ value then we get $q_{0}$ in the range $3\times 10^{-3}$ and $11\times 10^{-3}$ erg cm-3 s-1. For 335 Å it is $a=0.5$, $b=1.5$, and $q_{0}=4\times 10^{-3}$ erg cm-3 s-1. The fits within 20% of the minimum include $q_{0}$ values from $3\times 10^{-3}$ to $7\times 10^{-3}$ erg cm-3 s-1.
@@
Table 1: 211 and 355 Å best-fit comparisons
A recurring feature in these plots of model-to-data comparison metrics is the clear diagonal feature ranging from low $a$, high $b$ values on the top-left to high $a$, low $b$ values towards the bottom-right. This feature is consistent with other studies with EUV (Ugarte-Urra et al., 2019) and microwave data (Nita et al., 2023; Fleishman et al., 2025b; Kuznetsov et al., 2025). For the 8″ resolution 211 Å fit we found that this feature roughly corresponds to $b\approx 1.6-0.6a$ and for the 335 Å fits we get $b\approx 1.7-0.6a$. That our best fits for each wave band do not fall exactly on these lines is indicative of the coarseness of our ($a,b$) grid and uncertainties in our fitting of diagonal features. More specifically, if we assume a 0.25 uncertainty (half a ($a,b$) grid step) on both $a$ and $b$ we have $b$ intercepts of $1.61\pm 0.12$ and $1.70\pm 0.13$ and slopes of $0.56\pm 0.07$ and $0.59\pm 0.08$ for 211 Å and 335 Å, respectively, so the results for the two bands are the same within uncertainties.
We note that although the best $(a,b)$ values are different for the two wave bands, they both fall close to these best-fit lines. Table 1 compares the best model-to-data comparisons of the two wave bands. We see that for both wave bands the best ($a,b$) values fall within 10% of the minimum $\sum\delta^{2}$ for ($a,b$) values of the other wave band. This gives an idea, albeit somewhat qualitative, of our confidence in our fit values. The model-to-data comparisons for the other bands are sufficiently poor (see Sections III.2 and III.3) that we do not include them in this analysis.
@@
$B_{\mbox{\scriptsize avg}}\propto L^{-\gamma}$ (2)
(Klimchuk and Porter, 1995) then we have $\langle Q\rangle\propto L^{-\gamma a-b}$. Model fits should be similarly good when $\gamma a+b=$ constant; call it $b_{0}$. Good fits should therefore fall along a straight line in a plot of goodness of fit in the $(a,b)$ plane:
@@
Our results shown in Figure 15 suggest that $b_{0}\approx 1.6$ or 1.7 and $\gamma\approx 0.6$. As stated above, other analyses also show that similarly good fits fall along a straight line. Fleishman et al. (2025b) using GX Simulator to model optically thick gyroresonant microwave emission observed by the Siberian Solar Radio Telescope (SSRT) and Nobeyama Radioheliograph (NoRH) found b=2.41.0ab=2.4-1.0a at 5.7 GHz and b=2.20.7ab=2.2-0.7a at 17 GHz and a best solution at ($a,b$) = (0.6, 1.8). Kuznetsov et al. (2025) analyzed eight active regions in 6, 8, and 10 GHz with the Siberian Radioheliograph and found that $b_{0}$ was mostly between 1.6 and 3.2, with the distribution peaking at 2.5, and $\gamma$ ranged from 0.2 to 1.2, with a most probable value of 0.55. It is also interesting that Porter and Klimchuk (1995) concluded that $\langle Q\rangle\propto L^{-2}$ from a study of soft X-ray loops observed by Yohkoh, which can be compared to the $b_{0}$ values derived from GX Simulator.
@@
With this in mind, we fit the $B_{\mbox{\scriptsize avg}}$ vs $L$ dependence in our Magnetic Model 1, as shown in Figure 16, using a least-squares fit in log-log space. We derived $\gamma=0.8$. The plot also shows a line with a slope corresponding to $\gamma=0.6$ as suggested by the GX Simulator modeling. The value derived directly from the magnetic model is slightly higher than the value derived from $\sum\delta^{2}$ plots, but it is clear from the distribution of points in Fig. 16 that the distribution of the field lines in the model is more complex that a simple power-law, with noticeable substructure. We expect Equation 2 to be a general trend rather than an equation applying precisely to all parts of the active region. These results are, however, consistent with the idea that the diagonal feature in GX Simulator model-to-data comparison metrics is due to a dependence of $B_{\mbox{\scriptsize avg}}$ on $L$.
@@
Figure 16: Density distribution in the $(B_{\mathrm{avg}},\,L)$ plane for Magnetic Model 1 (2D histogram, equal log–log bins). Lines shown are an ordinary least-squares (OLS) regression to the entire set of points (solid blue line) and a line with a slope of $\gamma=0.6$, as derived from the GX Simulator fits (red dashed line). This line is forced to go through the center of mass of the magnetic model points (red circle).
@@
We should note that Fleishman et al. (2025b) and Kuznetsov et al. (2025) are looking at different active regions, which in addition to a difference in wave bands and a different model-to-data comparison metric could lead to differences in the properties of the zones of best fit. Kuznetsov et al. (2025) in particular find a range of values over the many active regions they modeled.
@@
However, the broad agreement of these results is very encouraging. This validates the power-law dependencies assumed in Equations 1 and 2. Had similarly good fits not fallen along a straight line, one or more of the dependencies would be invalidated.
IV.2 Transition-Region Emission
@@
An important issue in our modeling effort is that the emission from long features at the periphery of the active region is missing from the models. For this active region these are dominant features in the cooler bands, 131, and 171 Å, and significant contributors to other bands with peak contributions below about 1.6 MK (94 and 193 Å). We think this is due to an unrealistic distribution of the transition-region emission in these long loops. In this work and other modeling studies, the transition region is defined to be the locations in the loops where thermal conduction serves as a heating term rather than a cooling term (Vesecky et al., 1979). As discussed in Klimchuk et al. (2008), this is a precise, physically motivated definition that avoids ambiguities that arise from vague observational definitions based on temperature or altitude. As physically defined, the transition region extends from the top of the chromosphere to about 10% of a loop half-length ($0.05L$) (see Klimchuk et al., 2008; Cargill et al., 2022). An example of this is shown in Figure 17, which depicts the temperature, electron density, and normalized 171 Å intensity as a function of position of a 1D hydrodynamic model of a loop in equilibrium with steady uniform heating. The model was computed with the ARGOS code (Antiochos et al., 1999) and is broadly representative of static equilibrium loops with a wide range of lengths and heating rates. The vertical line shows the top of the transition region at $0.054L$ from the loop footpoints. This same spatial fraction applies approximately to other loops, but the temperature of the top of the transition region varies, always about 60% of the apex temperature.
@@
Figure 17: The temperature, electron density, and normalized 171 Å intensity of half of an equilibrium loop with uniform heating as modeled by ARGOS. The vertical line shows the location of the top of the transition region/base of the corona. It is about 5.4% of the full loop length, $L$, from the top of the chromosphere
@@
For relatively short loops it is sufficient to assume that the transition-region emission would be confined to a limited range approximated by a single voxel along the loop in our model cube. In our higher-resolution Model 1 each voxel is 1.4 Mm on a side. That would be 5% of a loop with $L=28$ Mm, within a factor of two or so of loop length for the core of an active region. However, based on Magnetic Model 2, the loops at the periphery of the active region likely have L=200700L=200-700 Mm, implying transition regions on the order of 10-50″ in the AIA images, much larger than the pixels immediately around the loop base and roughly consistent with the large cool loops we see in the data. Thus, in order to get a better fit to EUV images showing extended emission in loop legs, it will be necessary to account for the finite spatial extent of the transition region. Note that it is only the upper transition region, the upper third as determined by temperature, where this is important. The temperature scale lengths in the lower and even middle transition region are extremely short. The amount of transition-region emission predicted by EBTEL is reasonable (although see caveats about general transition-region modeling below), but because all of it is assigned to the footpoints by GX Simulator, the footpoints are too bright and the extended legs do not appear in model images.
@@
This explanation would explain Feature A, the fan-loop feature in Figure 11. It is not as clear if it accounts for Feature B in that plot, in which the emission appears to extend farther than $0.05L$ along a magnetic tube. We think it unlikely Feature A extends into the solar wind because we would expect the emission along open magnetic field lines to be fainter, as they are in coronal holes. It is possible the actual feature is on closed magnetic field lines that extend outside the volume we are modeling. However, that would still leave the question of why we would see emission on such a structure but not along the model field lines that are in Model 2.
@@
As discussed in Fleishman et al. (2025b), one possible way to obtain more information about the transition region is to combine analysis of microwave emission, which is sensitive to only the corona, with EUV emission, which includes both transition region and coronal components. This will help to separate out the emission from the two sources. Another approach would be to consider active regions from multiple points of view using AIA combined with either STEREO/EUVI (Solar TErrestrial RElations Observatory/Extreme UltraViolet Imager; Howard et al., 2008) or Solar Orbiter/EUI (Extreme Ultraviolet Imager; Rochus et al., 2020) near quadrature, so that lower-altitude transition-region emission could be separated out. However, for both of these approaches it would be important to consider the extended transition-region emission from the long loops discussed in this paper.
IV.3 Loop Expansion
We note an inconsistency in our models. Loops expand with height in the magnetic field extrapolation of GX Simulator, whereas the EBTEL simulations assume a constant loop cross section. Often, multiple voxels in a cut across the magnetic field map magnetically to a single transition-region voxel. Each of the coronal voxels is assigned a coronal DEM that corresponds to a companion transition-region DEM from the EBTEL simulation. The transition-region DEM appears one time in the active region model, but the coronal DEM appears multiple times. Consequently, these geometrical considerations suggest that the volume integrated coronal emission in the model is overestimated relative to the transition-region emission by a factor approximately equal to the area expansion factor of the loop.
However, this geometric effect is at least partially offset by the different thermodynamic properties of expanding and non-expanding loops. Cargill et al. (2022, Fig. 2) found that the ratio of coronal to transition-region emission per unit area is approximately independent of the expansion factor for a fixed heating rate. Therefore, for a given amount of transition-region emission, the volume integrated coronal emission should scale roughly with the expansion factor. This means that the thermodynamic effects missing from the EBTEL models (because they assume a constant cross section) fortuitously offset the geometric effects not accounted for in GX Simulator. In this sense, the current modeling is approximately correct. These effects must, of course, be investigated and verified more rigorously in future work. We note that EBTEL has the capability of simulating expanding loops, and this feature may need to be implemented in next-generation models. This is a significant undertaking as it adds a third dimension to the model lookup table.
We note also that Cargill et al. (2022) find that expansion can affect the ratio of the length of the transition region to $L$, so incorporating loop expansion may also affect the treatment of the location of the transition region in the model (see discussion in Section IV.2 above.)
IV.4 Further Considerations
@@
In addition to the treatment of the transition region and loop expansion there are other areas in which our modeling could be improved. As stated above, the details of the core magnetic field lines do not reflect the data well. This is likely because this active region had a rather weak magnetic field. It was selected because we wished to have a very quiet active region so as to avoid solar flares in the time-average images. Also, we wished to avoid sunspots because it is highly unlikely that sunspots obey the same heating scaling law that applies throughout most of the active region. For one thing, loops rooted in sunspot umbrae have much different emission properties from other active region loops; they are cooler and fainter (Tiwari et al., 2017). For another, many coronal heating models are based on the small-scale stressing of the field by photospheric convection, and convection is suppressed in sunspots. We therefore preferred to avoid this complication in our initial study, but of course sunspots must be included in an eventual complete model. A more typical active region, with stronger fields, may be easier to model magnetically. It would optimally make sense to have an iterative fit in which both the plasma and magnetic field parameters could be varied (see Fleishman et al. (2021a) for a case where fits to radio data were done using multiple magnetic models), but that is beyond the scope of this paper and our readily available computing power.
@@
Another area to consider is the assumptions going into the heating calculations in the lookup tables. Fleishman et al. (2025b) and Kuznetsov et al. (2025) considered the various lookup tables currently available in GX Simulator and, in analyses utilizing microwaves, concluded that the best fits for the diffuse coronal component come from assuming heating parameters based on higher frequency nanoflares than we used for this study or even constant heating. So, ultimately, heating with various nanoflare distributions should be considered for the EUV as well. Furthermore, the heating frequency may vary across the active region (Mondal et al., 2025), which should be incorporated into future models. Power-law forms for the nanoflare energy distribution are commonly assumed in the modeling community (e.g., Pauluhn and Solanki, 2007; Hood et al., 2016; Knizhnik et al., 2018) and are motivated by observed distributions (see review by Reale, 2014). However, the observed distributions are typically based on microflares and full-sized flares, and it is not obvious that the distributions can be extrapolated to smaller energies. Furthermore, the measured events are volume-integrated, whereas the nanoflares inputs to the EBTEL models are energy per unit cross-sectional area. Our multi-strand MHD simulations indicate that the energy per unit area actually obeys a log-normal distribution (Sow Mondal et al., 2025). The MHD simulations also indicate that the delay and energy of the nanoflares are uncorrelated (Sow Mondal et al., 2026), contrary to the assumption in the current GX Simulator lookup tables. These properties must be included when new lookup tables are computed.
@@
Although we concluded that it would not have helped in our case, it would be good to develop a way to estimate the loop lengths of field lines that exit the model box. At the moment the options are to either ignore the plasma in such lines or assume it is hydrostatic, neither of which is consistent with the modeling of the loops that close in the box. The calculation and manipulation of large volume NLFFF models needed to include more loops within the modeled box uses a lot of computer resources, so a way of estimating $L$ and $B_{\mbox{\scriptsize avg}}$ for large loops exiting smaller boxes would be helpful.
V Summary and Conclusions